# On the geometric phase and its role in the design of elastic topological materials

### v1

@@ Abstract
The geometric phase provides important mathematical insights to understand the occurrence and evolution of the dynamic response in a diverse spectrum of systems ranging from quantum to classical mechanics. While the concept of geometric phase, which is an additional phase factor occurring in dynamical systems, holds the same meaning across different fields of application, its use and interpretation can acquire important nuances specific to the system of interest. In recent years, the development of the concept of quantum topological materials and its extension to classical mechanical systems have renewed the interest in the study of the geometric phase. This study reviews the concept of geometric phase and discusses, by means of either established or original results, its role in the design of elastic materials. Concepts of differential geometry and topology are put forward to provide a theoretical understanding of the geometric phase and its connection to the physical properties of the system. Then, the concept of geometric phase is applied to different types of elastic waveguides to explain how either topologically trivial or non-trivial behavior can emerge based on a proper geometric design of the waveguide.
@@ 1 Introduction
The term geometric phase indicates a phase factor that is acquired by a dynamical system following an adiabatic perturbation of the system parameters at the end of which the system is brought back to the original conditions. The geometric phase is in addition to the usual dynamic phase accumulated by a harmonically oscillating system. While first recognized in the context of quantum mechanics [1], it was later observed also in other fields, such as optics [2], solid state physics [3], classical mechanics [4], and molecular physics [5, 6]. Over time, the concept of geometric phase has been extended also to nonadiabatic [7, 8] and nonlinear dissipative systems [9, 10, 11], its mathematical foundation has been firmly established [12, 13, 14], and it has found direct applications in optics [15, 16] and solid state physics [17, 18].
The geometric phase has proven to be a powerful parameter that provides mathematical insight into physical phenomena. The ubiquity and usefulness of the geometric phase across several disciplines of wave physics and dynamics arise from its rigorous mathematical description achievable via differential geometry and topology [14]. For example, the geometric phase concept elegantly describes the precession angle of a Foucault pendulum as a consequence of the curvature of the Earth, without having to explicitly solve the equations of motion [19]. Another example is the influence of the geometric phase on the rapidly developing field of elastic topological metamaterials (ETMs) [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Over the past decade, ETMs have opened new ways to control and manipulate the propagation of waves in elastic waveguides achieving unique behavior such as the propagation immune to back scattering produced by inhomogeneities and defects.
The concept of geometric phase has been discussed from a variety of perspectives, given its versatility and applicability to a multitude of applications. However, there is certainly a prevalence of discussions and applications within the quantum mechanics and photonics areas. In view of the rising interest in the use of the concept of geometric phase to study the dynamics of structures and materials, a cohesive discussion of its emergence and applications to the general area of elastic systems appears timely and necessary. While some discussions and results are present in the literature, their scattered and incomplete nature makes approaching this topic and understanding the state of the art quite challenging. As an example, most of the mathematical description of geometric phase is available in theoretical physics [14, 12, 13], while applications of geometric phase are briefly discussed in the context of topological metamaterials [31, 32, 33, 34, 22, 23, 26], and independent observations of geometric phases have been reported for elastic waves [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
In view of the above considerations, the present work aims at reviewing the concept of geometric phase in the context of elastic waves and at leveraging it to introduce the general topic of ETMs. The following discussion will also take advantage of notions of differential geometry and topology, but it will not require any prior knowledge.
@@
The article is organized as follows. Section 2 introduces the concept of geometric phase for systems described by a discrete set of variables. Section 3 reviews the formulation for elastic waveguides and provides examples of the geometric phase. Section 4 introduces some relevant concepts from differential geometry that are applied in Section 5 to explain the emergence of the geometric phase in elastic systems. Section 6 discusses the connection between the more theoretical mathematical ideas and the more practical dynamic behavior of elastic waveguides and ETMs. Section 7 summarizes the key points of this work.
@@ 2 Geometric phase in discrete systems
The geometric phase appears in a wide range of oscillating dynamical systems whose response is a function of a set of parameters, denoted in the following by the vector $\mathbf{R}$. Its appearance is well-studied in dynamical systems described by $n$ state variables, including discrete mechanical systems [4, 43]. Following the original formulation [1], the governing equation can be written as
$\mathcal{D}_{t}\mathbf{x}+\mathcal{L}(\mathbf{R})\mathbf{x}=0\;,$ (1)
where $\mathbf{x}$ is the vector of state variables, $t$ is time, $\mathcal{D}_{t}$ is a linear differential operator (typically $\mathcal{D}_{t}=d^{2}/dt^{2}$ in discrete mechanical systems, and $\mathcal{D}_{t}=\mathrm{i}d/dt$ in quantum mechanical systems), and $\mathcal{L}(\mathbf{R})$ is a matrix that depends on the parameters.
@@
where the explicit dependence of the eigenvalues $\lambda$ and the eigenvectors $\mathbf{X}$ on the parameters is highlighted. If $\mathbf{X}$ is an eigenvector, then also $\mathbf{X}e^{\mathrm{i}\gamma}$ is an eigenvector of the system, if $\gamma$ ranges between 00 and $2\pi$. In this sense, the eigenvector has an “ambiguity” in phase. If two eigenvectors at a parameter value $\mathbf{R}_{0}$ have the same eigenvalue and differ by more than just a phase term, they are said to be degenerate. The parameter value $\mathbf{R}_{0}$ is said to be a degeneracy. In a general scenario, spanning the parameter space does not give rise to degeneracies other than for a few isolated points. These selected values of the parameters are referred to as “accidental degeneracies”. However, if particular constraints (such as, for example, those arising from symmetries in the geometric configuration of the system) are present, then degeneracies will arise independently of the specific values of the parameters; or, equivalently, the degeneracies will not be lifted by simply choosing different values of the parameters (like in the case of accidental degeneracies). This latter class of degeneracies are said to be symmetry protected.
Suppose the parameters $\mathbf{R}(t)$ complete an adiabatic cycle in $T$ units of time. A cycle of the system parameters indicates that, in the parameter space, the point describing the properties of the system traces a closed loop $L$ (i.e., bringing the system back to the initial set of parameters). The adiabatic variation implies that the time-scale of the parameter variation is much larger than the difference in the time-scale of oscillations of the excited modes (i.e., $T\gg\mathrm{max}\left(\frac{1}{\lambda_{i}-\lambda_{j}}\right)$, where $i,j=1,\dots,n$).
@@
$\mathbf{x}(t)=\mathbf{X}(\mathbf{R}(t))\exp(\mathrm{i}\gamma(t))\exp\left(-\mathrm{i}\int_{0}^{t}\lambda(\mathbf{R}(t^{\prime}))\,\mathrm{d}t^{\prime}\right)\;.$ (3)
where $\gamma$ is the geometric phase that accounts for the phase ambiguity of the eigenvector. It is computed as [1]
@@
where $\mathbf{X}^{*}$ denotes the conjugate transposed. The derivation of Eq. (4) benefits from the mathematical discussions of Sec. 4 and it is postponed to Sec. 5. Equation (3) asserts that the state of a system at a time instant $t$ is determined by the mode of oscillation at the instantaneous parameter value $\mathbf{R}(t)$, the phase accumulated from oscillations, and the geometric phase.
@@
In this perspective, the geometric phase could be seen as an “additional memory” of the system in addition to the (dynamical) “phase memory” [41].
@@ 3 Geometric phases in elastic waves
The discussion of the occurrence and role of geometric phases in elastic continuous systems is somewhat scarce and scattered. To address this aspect and provide a more cohesive discussion of this concept from the perspective of elastic systems, we choose a prototypical system consisting in an isotropic quasi-one-dimensional (quasi-1D) waveguide subject to a cyclic and adiabatic variation of the cross section. By using this example, we will show how the geometric phase can arise [35]. For quasi-1D waveguides, an adiabatic parameter variation means that the length-scale of parameter variation is much larger than the inverse of the difference in wavenumbers of propagating waves. The analysis can be generalized to anisotropic and bulk materials [35, 41]. We discuss the geometric phase in an anisotropic bulk material in the supplementary material (Sec. 1)).
There are two phase components available in elastic waves: the overall phase accumulated by a wave packet while traveling through the medium and the polarization angle, which is relevant in waveguides supporting polarized waves. The polarization angle specifies the “orientation” of the displacement field at a transverse plane of the waveguide, in a manner that will be further clarified in Sec. 3.2. A geometric phase can modify either one of the two phase components.
The mathematical description of the geometric phase proceeds from the analysis of the Navier-Lamé (NL) equations of elastodynamics with traction-free boundary conditions. A quasi-1D waveguide whose longitudinal axis coincides with the $z$ axis is considered. The geometric properties of the cross section at $z$ are defined by the parameters $\mathbf{R}(z)$. For this waveguide, the NL equations and traction-free boundary conditions read [46]
@@
respectively, where $\rho$ is the density, $\mathbf{u}$ is the displacement field, $\lambda$ and $\mu$ are the Lamé constants, $\sigma$ is the stress tensor, and $\hat{n}$ is the outward unit normal at the surface. The stress tensor for an isotropic elastic solid is [47] $\sigma=C:\frac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{\intercal})$, where $C$ is the fourth order stiffness tensor.
If the parameters $\mathbf{R}$ are constant, the governing equations admit plane wave solutions of the form $\mathbf{u}(x,y,z,t)=\mathbf{U}(x,y)e^{\mathrm{i}(kz-\omega t)}$, where $\omega$ is the angular frequency, $k$ is the wavenumber, and $\mathbf{U}(x,y)$ is the amplitude of the displacement function. The term in the exponent $\phi_{d}=kz-\omega t$ is the dynamical phase. These plane wave solutions are the guided modes of the waveguide that physically represent displacement field disturbances propagating along the longitudinal $z$ direction without changes in their displacement amplitudes $\mathbf{U}(x,y)$. Guided modes are determined by substituting the plane wave ansatz into Eqs. (6), (7). This results in
@@
where $\mathcal{L}(k,\mathbf{R})$ and $\mathcal{B}(\mathbf{R})$ are linear differential operators. Equations (8), (9) define an eigenvalue problem at each value of $k$. The eigenvalues define the so-called dispersion relations that link the angular frequency to the wavenumber for modes propagating in the waveguide. The associated displacement amplitude functions define the particle displacement fields associated with the guided modes. Further, the dependence of the dispersion relation and of the displacement amplitude functions on the parameter $\mathbf{R}$ is explicitly written in Eqs. (8), (9). If two guided modes at a parameter value $\mathbf{R}_{0}$ have identical eigenvalues for all values of $k$ (i.e. identical dispersion relations), we will call the associated guided modes as “fully degenerate”. The parameter value $\mathbf{R}_{0}$ will be called a “full degeneracy”. This choice of terminology is used to distinguish it from degeneracy in an eigenvalue problem (Sec. 2) that does not depend on a parameter such as $k$.
Now consider a semi-infinite waveguide with the geometric properties $\mathbf{R}$ varying adiabatically along the positive $z$ direction. Further, assume that the properties $\mathbf{R}(z)$ vary in a cycle between $z=0$ and $z=Z$, i.e., the properties at $z=0$ are identical to those at $z=Z$, written as $\mathbf{R}(0)=\mathbf{R}(Z)$. Adiabatic variation along $z$ implies that disturbances in the waveguide propagate according to the WKB approximation[35, 48]. In particular, if the waveguide is harmonically excited with angular frequency $\omega$ by the displacement amplitude $\mathbf{U}(x,y;\mathbf{R}(0))$ at $z=0$, the steady state displacement field at $z$ is described by: (i) a displacement amplitude $\mathbf{U}(x,y;\mathbf{R}(z))$, (ii) an amplitude factor $a(z)$ computed by the conservation of an adiabatic invariant, (iii) the usual dynamical component of phase $\phi_{d}$ computed as
@@
and (iv) the geometric phase. The computation of the geometric phase is deferred to Sec. 5 as it can benefit from concepts of topology and differential geometry.

### v3

# On the role of geometric phase in the dynamics of elastic waveguides

@@ Abstract
The geometric phase provides important mathematical insights to understand the fundamental nature and evolution of the dynamic response in a wide spectrum of systems ranging from quantum to classical mechanics. While the concept of geometric phase, which is an additional phase factor occurring in dynamical systems, holds the same meaning across different fields of application, its use and interpretation can acquire important nuances specific to the system of interest. In recent years, the development of quantum topological materials and its extension to classical mechanical systems have renewed the interest in the concept of geometric phase. This review revisits the concept of geometric phase and discusses, by means of either established or original results, its critical role in the design and dynamic behavior of elastic waveguides. Concepts of differential geometry and topology are put forward to provide a theoretical understanding of the geometric phase and its connection to the physical properties of the system. Then, the concept of geometric phase is applied to different types of elastic waveguides to explain how either topologically trivial or non-trivial behavior can emerge based on the geometric features of the waveguide.
@@ 1 Introduction
The term geometric phase indicates a phase factor acquired by a dynamical system following an adiabatic perturbation of the system parameters at the end of which the system is brought back to the original conditions. The geometric phase is in addition to the usual dynamical phase accumulated by a harmonically oscillating system. While first recognized in the context of quantum mechanics [1], it was later observed also in other fields, such as optics [2, 3, 4], solid state physics [5, 6, 7], classical mechanics [8], and molecular physics [9, 10]. The concept of geometric phase has been extended also to nonadiabatic [11, 12] and nonlinear dissipative systems [13, 14, 15], its mathematical foundation has been firmly established [16, 17, 18].
The geometric phase has proven to be a powerful concept to provide deep mathematical insight into certain aspects of dynamic phenomena. The ubiquity and usefulness of the geometric phase across several disciplines of wave physics and dynamics arise from its rigorous mathematical description and its interpretation via differential geometry and topology [18]. For example, the geometric phase concept can describe the precession angle of a Foucault pendulum [19], provide an approach to sense the spatial distribution of scatterers in acoustic scattering problems [20, 21], model parameter variations required to realize non-Abelian physics [22, 23], and explain mathematical foundations and design principles underlying elastic topological metamaterials [24, 25, 26, 27, 28, 29, 30]. In this article, we will focus on the latter application as a way to provide a concrete example to illustrate the role of the geometric phase in the design of dynamical systems.
In recent years, topological metamaterials have gathered the interest of the physics and engineering communities. While the concept of topological metamaterial has found application in many areas of physics, in the following we focus on elastic topological metamaterials (ETMs) as a practical example to discuss the role of the geometric phase. Certainly many of these concepts are general and could be readily extended, for example, to photonics and phononics applications. ETMs can be seen as a class of waveguides whose dynamic response can be understood, and even engineered, by using concepts of differential geometry and topology. As an example, over the past decade, ETMs have opened new ways to control and manipulate the propagation of stress waves. A very distinctive trademark of topological materials is the ability to support elastic waves (typically on boundaries or interfaces characterized by a topological transition) that are immune to back scattering from inhomogeneities and defects [31, 32, 33]. The ETM field has grown extremely rapidly during the last decade also capitalizing on the many decades of research on (quantum) topological insulators [34, 24, 26], which first exploited the concept of topologically distinct classes of materials. The extension of these concepts to classical (i.e. non quantum mechanical) platforms, such as photonic [35, 36] and acoustic systems [30, 37], also represented another key milestone to illustrate the feasibility of topological concepts for ETMs. Given that this review focuses on the role of geometric phase in the description of the dynamic behavior of elastic systems and uses ETMs only as an example of a potential application, a detailed discussion of ETMs is omitted. Nevertheless, the interested reader can refer to the following papers [38, 31, 32, 33, 39] and reviews [37, 28, 29, 40, 30, 41, 42] as a starting point to explore the ETM field more in detail.
In view of the rising interest in using the geometric phase to study the dynamics of structures and materials, a cohesive discussion of its features and potential applications appears timely and necessary. This is not to say that the literature does not offer discussions on the geometric phase, but these discussions are often geared towards a quantum mechanical or theoretical physics audience [18, 16, 17, 24]. The few discussions focused on applied mechanics tend to be scattered and limited to a narrow set of applications [30, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 20, 21, 54], therefore making approaching this topic and understanding the state of the art quite challenging.
In view of the above considerations, the present article aims to review the concept of geometric phase in the context of elastic waveguides and to leverage it to explain some fundamental ideas in ETMs.
@@
The review is organized as follows. Section 2 introduces the concept of geometric phase for systems described by a discrete set of variables. Section 3 reviews the formulation for elastic waveguides and provides examples of geometric phase. Section 4 introduces some relevant concepts from differential geometry that are applied in Section 5 to explain the emergence of the geometric phase in elastic systems. Section 6 discusses the connection between the more theoretical mathematical ideas and the more practical dynamic behavior of elastic waveguides and ETMs. Section 7 summarizes the key points of this work.
@@ 2 Geometric phase in discrete systems
The geometric phase appears in a wide range of oscillatory dynamical systems whose response depends on externally controlled time-varying parameters. Its appearance is well-studied in dynamical systems described by $n$ state variables, including discrete mechanical systems [8, 51]. Following the original formulation [1], the governing equation can be written as
$\mathcal{D}_{t}\mathbf{x}+\mathcal{L}(\mathbf{R}(t))\mathbf{x}=0\;,$ (1)
where $\mathbf{x}$ is the vector of state variables, $\mathbf{R}(t)$ is the vector of controllable parameters, $t$ is time, $\mathcal{D}_{t}$ is a linear differential operator (typically $\mathcal{D}_{t}=d^{2}/dt^{2}$ in discrete mechanical systems, and $\mathcal{D}_{t}=\mathrm{i}d/dt$ in quantum mechanical systems), and $\mathcal{L}(\mathbf{R})$ is a matrix that depends on the parameters.
@@
where the explicit dependence of the eigenvalues $\lambda$ and the eigenvectors $\mathbf{X}$ on the parameters is highlighted. Note that here $\mathbf{R}$ is constant by assumption If $\mathbf{X}$ is an eigenvector, then also $\mathbf{X}e^{\mathrm{i}\gamma}$ is an eigenvector of the system, if $\gamma$ ranges between 00 and $2\pi$. In this sense, the eigenvector has an “ambiguity” in phase. If two eigenvectors at a parameter value $\mathbf{R}_{0}$ have the same eigenvalue and differ by more than just a phase term, they are said to be degenerate. The parameter value $\mathbf{R}_{0}$ is said to be a degeneracy. In a general scenario, at least three parameters must be independently controlled to result in a degeneracy [1]. That is, degeneracies have codimension 3. These selected values of the parameters are referred to as “accidental degeneracies”. However, if particular constraints (such as, for example, those arising from symmetries in the geometric configuration of the system) are present, then degeneracies will arise independently of the specific values of the parameters; or, equivalently, the degeneracies will not be lifted by simply choosing different values of the parameters (as long as the constraints are not violated). This latter class of degeneracies is said to be symmetry protected.
Suppose the parameters $\mathbf{R}(t)$ complete an adiabatic cycle in $T$ units of time. A cycle of the system parameters indicates that, in the parameter space, the point describing the properties of the system traces a closed loop $L$ (i.e., bringing the system back to the initial set of parameters). The adiabatic variation dictates that the rate of change of parameters is “sufficiently slow”, hence implying that the time-scale of the parameters variation is much larger than the difference in the time-scale of oscillations of the excited modes (i.e., $T\gg\mathrm{max}\left(\frac{1}{\lambda_{i}-\lambda_{j}}\right)$, where $i,j=1,\dots,n$). Note that, in this context, the term “adiabatic” is used in its quantum mechanical sense (of a slowly-varying quantity). Therefore, it should not be interpreted in its thermodynamics sense, which instead indicates a “fast enough” process that prevents heat transfer between a system and its surroundings.
@@
$\mathbf{x}(t)=\mathbf{X}(\mathbf{R}(t))\exp(\mathrm{i}\gamma(t))\exp\left(-\mathrm{i}\int_{0}^{t}\lambda(\mathbf{R}(t^{\prime}))\,\mathrm{d}t^{\prime}\right)\;,$ (3)
where $\int_{0}^{t}\lambda(\mathbf{R}(t^{\prime}))\,\mathrm{d}t^{\prime}$ is referred to as the dynamical phase [16] and $\gamma$ is referred to as the geometric phase. In other words, the state of a system at a time instant $t$ is determined by the mode of oscillation at the instantaneous parameter value $\mathbf{R}(t)$, the dynamical phase, and the geometric phase. The dynamical phase describes the oscillatory nature of the system, for example, the phase acquired by a harmonic oscillator with a time-varying spring stiffness. If the parameters were constant, the dynamical phase would reduce to the familiar expression $\exp\left(-\mathrm{i}\lambda(\mathbf{R})t\right)$. The dynamical phase depends on the path traced in parameter space and the rate of parameter variation.
The geometric phase accounts for the phase ambiguity of the eigenvector. It is computed as [1]
@@
where $\mathbf{X}^{*}$ denotes the conjugate transposed. The derivation of Eq. (4) benefits from the mathematical discussions of Sec. 4 and it is postponed to Sec. 5. According to Eq. (4), the geometric phase depends only on the path traced in parameter space, not on the rate of parameter variation, in contrast to the dynamical phase.
@@
The role of the geometric phase can be intuitively understood as being an intrinsic memory of the system [49]. The phase acquired by an oscillatory system is, loosely speaking, the oscillator’s “memory” because it records the oscillator’s parameter variations over a given period of time (or, equivalently, the regions of the parameter space that the system has visited at a given time). This phase “memory” has two parts: the dynamical phase and the geometric phase. While the dynamical phase is the “memory” component familiar in classical (and quantum) physics, the geometric phase is an unexpected “additional memory”.
@@ 3 Geometric phases in elastic waves
The discussion of the occurrence and role of geometric phases in elastic continuous systems is somewhat scarce and scattered. To address this aspect and provide a more focused discussion of this concept from the perspective of elastic systems, we choose a prototypical system consisting in an isotropic quasi-one-dimensional (quasi-1D) waveguide subject to a cyclic and adiabatic variation of the cross section’s geometry. By using this example, we will illustrate the type of mechanisms leading to the accumulation of geometric phase [43]. For quasi-1D waveguides, an adiabatic parameter variation means that the length-scale over which the parameters vary is much larger than the inverse of the difference in wavenumbers of propagating waves. The analysis can be generalized to include anisotropic and bulk materials [43, 49]. The geometric phase in the context of anisotropic bulk materials is discussed in supplementary material (Sec. 1))
There are two phase components available in elastic waves: the overall phase accumulated by a wave packet while traveling through the medium and the polarization angle, which is relevant for waveguides supporting polarized waves. The polarization angle specifies the “orientation” of the displacement field at a transverse plane of the waveguide, in a manner that will be further clarified in Sec. 3.2. A geometric phase can modify either one of the two phase components.
The mathematical description of the geometric phase proceeds from the analysis of the Navier-Lamé (NL) equations of elastodynamics with traction-free boundary conditions. A quasi-1D waveguide whose longitudinal axis coincides with the $z$ axis is considered. The geometric properties of the cross section at $z$ are defined by the parameters $\mathbf{R}(z)$. For this waveguide, the NL equations and traction-free boundary conditions are [56]
@@
where $\rho$ is the density, $\mathbf{u}$ is the displacement field, $\lambda$ and $\mu$ are the Lamé constants, $\sigma$ is the stress tensor, and $\hat{n}$ is the outward unit normal at the surface. The stress tensor for an isotropic linear elastic solid is [57] $\sigma=C:\frac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{\intercal})$, where $C$ is the fourth order stiffness tensor.
If the parameters $\mathbf{R}$ are constant, the governing equations admit plane wave solutions of the form $\mathbf{u}(x,y,z,t)=\mathbf{U}(x,y)e^{\mathrm{i}(kz-\omega t)}$, where $\omega$ is the angular frequency, $k$ is the wavenumber, and $\mathbf{U}(x,y)$ is the displacement field of a cross section. The term in the exponent $\phi_{d}=kz-\omega t$ is the dynamical phase. These plane wave solutions are the guided modes of the waveguide that physically represent displacement field disturbances propagating along the longitudinal $z$ direction without changes in their cross-sectional displacement fields $\mathbf{U}(x,y)$. Guided modes are determined by substituting the plane wave ansatz into Eqs. (6), (7). This results in
@@
where $\mathcal{L}(k,\mathbf{R})$ and $\mathcal{B}(\mathbf{R})$ are linear differential operators. Equations (8), (9) define an eigenvalue problem at each value of $k$. The eigenvalues define the so-called dispersion relations that link the angular frequency to the wavenumber of modes propagating in the waveguide. The eigenfunctions $\mathbf{U}(x,y;k,\mathbf{R})$ define the cross-sectional particle displacement fields associated with the guided modes. Further, the dependence of the dispersion relations and of the eigenfunctions on the parameter $\mathbf{R}$ is explicitly indicated in Eqs. (8), (9). If two guided modes at a parameter value $\mathbf{R}_{0}$ have identical eigenvalues for all values of $k$ (i.e. identical dispersion relations), the associated guided modes are said to be “fully degenerate”. The parameter value $\mathbf{R}_{0}$ will be called a “full degeneracy”. This choice of terminology is used to distinguish it from degeneracy in an eigenvalue problem (Sec. 2) that does not depend on a parameter such as $k$.
The general setup to study the geometric phase in the remainder of the paper considers a waveguide of length $Z$ with the geometric properties $\mathbf{R}$ varying adiabatically along the positive $z$ direction. Further, assume that the properties $\mathbf{R}(z)$ vary in a cycle between $z=0$ and $z=Z$, i.e., the properties at $z=0$ are identical to those at $z=Z$, written as $\mathbf{R}(0)=\mathbf{R}(Z)$. If the waveguide is harmonically excited with angular frequency $\omega$ by the guided mode eigenfunction $\mathbf{U}(x,y;\mathbf{R}(0))$ at $z=0$, the steady state displacement field at $z$ is described by: (i) a guided mode eigenfunction $\mathbf{U}(x,y;\mathbf{R}(z))$, (ii) an amplitude factor $a(z)$ computed by the conservation of an adiabatic invariant, (iii) the usual dynamical component of phase $\phi_{d}$ computed as
@@
and (iv) the geometric phase [43, 58]. The computation of the geometric phase is deferred to Sec. 5 as it can benefit from concepts of topology and differential geometry that will be introduced in Sec. 4.